X2-X-6=0 then X is
Answer by stanbon(75887)

X^2-X-6=0 then X is
------
Factor to get:
(x-3)(x+2) = 0
Since the product is zero one of the factors must be zero.
x-3 = 0 or x+2 = 0
x = 3 or x = -2
=====================
Cheers,
Stan H.